# Fairness and Bias in Multimodal : A Survey

### v1

@@ 5.3.4 Other strategies
Oversampling under-represented groups in the data has been advocated by Buolamwini and Gebru (2018). There are generation detoxifying methods with the potential to reduce bias Gehman et al. (2020). These include data-based further pretraining and decoding-based generation. Similarly, Cabello et al. (2023) showed that continued pretraining on gender-neutral data improves fairness by reducing group disparities in some language-vision tasks. In another work, Yan et al. (2020) used adversarial learning. They added a discriminator to jointly learn with the predictor for the sensitive attributes. Meanwhile, Zayed et al. (2024) addressed fairness by pruning in LLM s while Guo et al. (2022) introduced auto-bias, by directly probing the biases in pretrained models through prompts. Liang et al. (2021) introduced the Autoregressive Iterative Nullspace Projection (A-INLP)method to carry out post-hoc debiasing on LLM s. Additionally, there are Self-Debias Schick et al. (2021), Hard-Debias Bolukbasi et al. (2016), SentenceDebias Liang et al. (2020) and Dropout methods Webster et al. (2020); Meade et al. (2022)
@@ 6 Conclusion
Fairness and bias are very important considerations in multimodal AI. In this work, we presented the challenges of fairness and bias in multimodal data, LMM s, and LLM s, defining what both terms mean within the scope of this survey, while acknowledging other definitions in the literature. We discussed the concepts of fairness and bias from the perspective of the Social Science and showed the distribution of scientific publications across 17 publishers, which reveals the gap of study in large multimodal AI compared to LLM s, which this work contributes to filling. Our discussion around the methods to measure fairness and bias, datasets for evaluation, and debiasing strategies will provide researchers and other stakeholders with insight on how to approach these issues. For future work, it will be worthwhile to re-evaluate the progress made with the metrics and tools for quantifying and mitigating fairness and bias concerns, respectively.
@@ Acknowledgements
This work is funded by the EU HumaneAI Project. The work is partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation and counterpart funding from Luleå University of Technology (LTU).

### v2

@@ 5.2.13 Other strategies
Furthermore, there are generation detoxifying methods with the potential to reduce bias Gehman et al. (2020). These include the earlier-mentioned continued pretraining and decoding-based generation. Buolamwini and Gebru (2018) advocated oversampling under-represented groups in data to mitigate bias. Zayed et al. (2024) addressed fairness by pruning in LLM s while Guo et al. (2022) introduced auto-bias, by directly probing the biases in pretrained models through prompts. Liang et al. (2021b) introduced the Autoregressive Iterative Nullspace Projection (A-INLP)method to carry out post-hoc debiasing on LLM s. Additionally, there are Self-Debias Schick et al. (2021), Hard-Debias Bolukbasi et al. (2016), SentenceDebias Liang et al. (2020) and Dropout methods Webster et al. (2020); Meade et al. (2022).
@@ 6 Conclusion
Fairness and bias are very important considerations in multimodal AI. In this work, we presented the challenges of fairness and bias in multimodal data, LMM s, and LLM s, defining what both terms mean within the scope of this survey, while acknowledging other definitions in the literature. We discussed the concepts of fairness and bias from the perspective of the Social Science and showed the distribution of scientific publications across many publishers, which reveals the gap in the study of large multimodal AI compared to LLM s, which this work contributes to filling. Our discussions around the methods to measure fairness and bias, datasets for evaluation, and debiasing strategies will provide researchers and other stakeholders with insight on how to approach these issues.
For future work, it will be worthwhile to re-evaluate the progress made with the metrics and tools for quantifying and mitigating the challenges of fairness and bias because despite the positive effects of debiasing in AI, researchers like West et al. (2019) argue that such research ought to do more than technical debiasing and include the social analysis of the use of such AI, as this will account more for the impact of bias overall Weinberg (2022). Even Sami et al. (2023) showed with the example of DALLE-2 that current efforts still have their limitations. In their work on the LMM DALLE-2, they revealed that, despite the guardrails for the model by OpenAI, it generated 40 more images of women for a stereotypical female-dominant administrative task, in clear gender bias, when prompted.
@@ Acknowledgements
This work is supported by the European Commission-funded project "Humane AI: Toward AI Systems That Augment and Empower Humans by Understanding Us, Our Society and the World Around Us." The authors wish to thank our colleagues at the Responsible Artificial Intelligence Group at Umeå University. The work is partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation and counterpart funding from Luleå University of Technology (LTU).
Appendix A Appendix
A.1 Google Scholar Papers
A.1.1 Multimodal444scholar.google.com/scholar?hl=en&as_sdt=0%2C5&q=Fairness+ and+Bias+in+Large+Multimodal+Models&btnG=
IEEE
: Wang et al. (2020), Niu et al. (2021a), Joshi et al. (2021), Wang et al. (2022c), Seth et al. (2023), Karkkainen and Joo (2021), Peng et al. (2022), Cho et al. (2023), Wang et al. (2022b), Niu et al. (2021b), Hirota et al. (2022b)
Springer
: Xu et al. (2020), Peña et al. (2023), Stoychev and Gunes (2022), Georgopoulos et al. (2021), Cai et al. (2024), Alwahaby et al. (2022), Kehrenberg et al. (2020)
ACM
: Booth et al. (2021), Yan et al. (2020), Wolfe and Caliskan (2022a), Goyal et al. (2022b), Cheong et al. (2024), Wolfe and Caliskan (2022b), Liang et al. (2022a), Yao et al. (2023), Hirota et al. (2022a), Yee et al. (2021), Rahman et al. (2020), Pessach and Shmueli (2022), Xu et al. (2024), Tang et al. (2021)
arXiv
: Luo et al. (2024), Zhou et al. (2021), Wang et al. (2024), Birhane et al. (2021), Friedrich et al. (2023), Goyal et al. (2022a), Zhang et al. (2024b), Rana and Jha (2022), Chen et al. (2024), Soares and Angelov (2019), Zong et al. (2022)
PubMed
: Liang et al. (2021a)
MDPI
: Ferrara (2023a), Nazi and Peng (2024)
ACL
: Wang et al. (2021), Srinivasan and Bisk (2022), Li et al. (2021), Zhang et al. (2024a), Ross et al. (2021)
Nature
: Meng et al. (2022a), Röösli et al. (2022), Fei et al. (2022)
PLoS
: Jiang et al. (2024b)
SPIE
: Drukker et al. (2023)
De Gruyter
: Jenks (2024)
NeurIPS
: Gadre et al. (2024), Liang et al. (2022c), Koh et al. (2024), Lee et al. (2023)
Elsevier
: Rahate et al. (2022), Serna et al. (2022), Kumari and Ekbal (2021)
PKP
: Jaiswal and Provost (2020)
MLRP
: Kiros et al. (2014), Zhang et al. (2022a)
Wiley
: Ntoutsi et al. (2020)
Sage
: Tay et al. (2022)
A.1.2 Language666scholar.google.com/scholar?hl=en&as_sdt=0%2C5&q= Fairness+and+Bias+in+Large+Language+Models&btnG=
NeurIPS
: Ma et al. (2023), Vig et al. (2020), Kojima et al. (2022), Solaiman and Dennison (2021), Wick et al. (2019), Tan and Celis (2019), Ouyang et al. (2022), Schaeffer et al. (2024), Meng et al. (2022b), Brown et al. (2020), Yu et al. (2023)
ACM
: Bender et al. (2021), Chang et al. (2024), Dhamala et al. (2021), Weidinger et al. (2022), Mehrabi et al. (2021), Garg et al. (2019), Ganguli et al. (2022), Min et al. (2023), Guo and Caliskan (2021), Zhang et al. (2020), Dodge et al. (2019), Navigli et al. (2023), Kotek et al. (2023)
PKP
: Pryzant et al. (2020)
MIT Press
: Ziems et al. (2024), Jiang et al. (2021), Schick et al. (2021), Raza et al. (2024)
Springer
: Freiberger and Buchmann (2024), Meyer et al. (2023), Teubner et al. (2023), Hou et al. (2024), Lu et al. (2020)
MLRP
: Liang et al. (2021b), Biderman et al. (2023), Aher et al. (2023), Lehman et al. (2023), Liang et al. (2021b)
JMLR
: Chung et al. (2024), Chowdhery et al. (2023)
Cambridge
: Argyle et al. (2023)
arXiv
: Ferrara (2023b), Wang et al. (2023a), Tamkin et al. (2021), Chen et al. (2021), Liang et al. (2022b), Weidinger et al. (2021), bench authors (2023), Rae et al. (2021), Eloundou et al. (2023), Serapio-García et al. (2023), Qi et al. (2023), Smith et al. (2022b), Huang et al. (2023), Liu et al. (2023a), Liu et al. (2023b), Hu et al. (2021), Zhang et al. (2022b), Sinha et al. (2021), Thoppilan et al. (2022), Menick et al. (2022), Doan et al. (2024), weidinger2019ethical
Elsevier
: Kasneci et al. (2023), Liu et al. (2023c), Ray (2023)
ACL
: Welbl et al. (2021), Nadeem et al. (2021), Gehman et al. (2020), Delobelle et al. (2020), Sap et al. (2020), Park et al. (2018), Perez et al. (2022), Blodgett et al. (2020), Khanuja et al. (2023), Agrawal et al. (2022), Sun et al. (2019), Liang et al. (2020), Dinan et al. (2020), Blodgett et al. (2021), Chen and Mueller (2024), Gao et al. (2020b), Kurita et al. (2019), Yoo et al. (2021), Lin et al. (2022), Hutchinson et al. (2020), Ramesh et al. (2023), Smith et al. (2022a), Ruder et al. (2022)
Nature
: Schramowski et al. (2022), Clusmann et al. (2023), Hager et al. (2024), Meskó and Topol (2023), Thirunavukarasu et al. (2023)
PLoS
: Mozafari et al. (2020)
Wiley
: Hovy and Prabhumoye (2021)
MDPI
: Garrido-Muñoz et al. (2021)
Preprints
: Kumar et al. (2024)
PubMed
: Karabacak and Margetis (2023)
Academic Pinnacle
: Desai et al. (2023)
A.2 Web of Science (WoS)Papers
A.2.1 Multimodal [^8] https://www.webofscience.com/wos/woscc/summary/889a9d49-d906-408c-91b6-ffefdff1880d-fed7fb8c/relevance/1
IEEE
: Seth et al. (2023), Porgali et al. (2023), Sami et al. (2023), Petryk et al. (2022)
ACM
: Edenberg and Wood (2023), Mandal et al. (2023b), Alam (2022)
PKP
: Lui et al. (2024)
A.2.2 Language101010www.webofscience.com/wos/woscc/summary/79cd811f-beb4-40f2-844f-e0ca9edf1fe2-fed7f83e/relevance/1
IEEE
: Malic et al. (2023), Li et al. (2024), Morales et al. (2023), Da et al. (2024), Jeong et al. (2024)
Elsevier
: Nazir et al. (2024), Fan et al. (2020)
ACM
: Dacon and Liu (2021), Ovalle et al. (2023), Halevy et al. (2021), Mei et al. (2023), Salinas et al. (2023), Dhamala et al. (2021), Gadiraju et al. (2023), Saxena et al. (2024), Jamil (2024), Zerveas et al. (2022), Wright et al. (2020), Giner-Miguelez et al. (2023), Raj et al. (2023)
Springer
: Dolci et al. (2023), Delobelle and Berendt (2022), Schröder et al. (2023), Leteno et al. (2023), Yang et al. (2022),
NeurIPS
: Ma et al. (2023)
Nature
: Haltaufderheide and Ranisch (2024)
MDPI
: Bevara et al. (2024), Piñeiro-Martín et al. (2023), da Silva et al. (2021)
MLRP
: Liang et al. (2021c)
Now
: Xue et al. (2023)
Wiley
: Lee et al. (2024), Hao et al. (2024)
ACL
: Kumar et al. (2023), Lauscher et al. (2021), Sheng et al. (2021), Fatemi et al. (2023), Guo et al. (2022), Vanmassenhove et al. (2018), Huang et al. (2020a), Felkner et al. (2023), Talat et al. (2022), Escudé Font and Costa-jussà (2019)
AIAF
: Van der Wal et al. (2024)